# Expanding the Chaos: Neural Operator for Stochastic (Partial) Differential Equations

### v1

@@ 5.1 Approximation of SPDE Solutions
We present learning outcomes in Table 2 and Figure 3, following the qualitative protocol used in NSPDE (salvi2022neural). $\mathcal{F}$-SPDENO attains the lowest error among all competitors in Table 2. Compared to the strongest baseline NSPDE, the error is reduced from $0.040$ to $0.037$ on $W\mapsto X$ (about $7.5\%$ relative improvement), and more substantially from $0.047$ to $0.031$ on $(X_{0},W)\mapsto X$ (about $34\%$ relative improvement). This indicates that our model better captures the joint effect of the random forcing and varying initial conditions. We also observe that deterministic operator learners such as FNO already perform competitively on $W\mapsto X$, but degrade notably when the initial condition varies (from $0.051$ to $0.073$). By contrast, $\mathcal{F}$-SPDENO consistently improves over FNO in both settings (from $0.051$ to $0.037$ and from $0.073$ to $0.031$), showing that enriching FNO with WCE-based noise features is crucial for handling stochastic SPDEs. In contrast to iterative inference methods like NSPDE, our framework first uses WCE to convert the stochastic forcing into static features. This enables a single FNO to learn the global solution operator in a one-shot forward pass, while still achieving superior accuracy.
@@ 5.2 Approximation on SDE Solutions
We now turn our attention to applying our model to SDEs. As we have mentioned, one of the key advantages of the NO method is that it enables one-shot evaluation. This naturally makes SDENO a one-step sampler for diffusion generative models (ho2020denoising; song2020score). In this experiment, we apply our $\mathcal{U}$-SDENO on pretrained DDPM models on CIFAR10 and CelebA-HQ  [^2] Model id :google/ddpm-cifar10-32 and google/ddpm-celebahq-256.. Specifically, the canonical DDPM reverse SDE takes the form
$dX_{t}\;=\;\Bigl[-\tfrac{1}{2}\,\beta(t)\,X_{t}\;-\;\beta(t)\,\nabla_{x}\log p_{t}(X_{t})\Bigr]\,dt\;+\;\sqrt{\beta(t)}\,d\overline{W}_{t}.$ (8)
At each reverse step $t$, we sample $\widehat{X}_{t}$ from pretrained models, and compute the DDPM posterior mean $\mu_{t}$ and variance $\sigma_{t}^{2}$ from the fixed noise schedule $\{\beta_{s}\}_{s=1}^{T}$, and then draws $X_{t-1}=\mu_{t}+\sigma_{t}z_{t}$ with $z_{t}\sim\mathcal{N}(0,I)$ (no noise at t=0t{=}0). Operationally, $z_{t}$ realizes the reverse-time increment $d\overline{W}_{t}\approx\sqrt{\Delta t}\,z_{t}$. We sample $d\overline{W}$ trajectories followed by the previous work in (zheng2023fast). Figure 4 shows the generated images of $\mathcal{U}$-SDENO at the last prediction step compared with the pretrained DDPM outputs on CIFAR-10, and unconditional samples on CelebA-HQ. One can see that $\mathcal{U}$-SDENO reconstructs the images with high fidelity and produces unconditional samples of comparable quality. Unlike standard diffusion models, which learn the score fields and integrate the reverse SDE step by step, SDENO learns the entire solution operator that maps a full noise trajectory $d\overline{W}_{1:T}$ directly to the final sample in a single forward pass. At sampling time, we only need to draw a Gaussian noise trajectory and evaluate SDENO once, without repeatedly invoking the large UNet denoiser. The goal of this experiment is therefore not to compete with specialized one-step diffusion schemes, e.g., consistency models (song2023consistency), but to demonstrate that the same WCE-based NO framework can also handle high-dimensional SDEs in the image domain.d
@@
In this experiment, we show SDENO also possesses strong interpolation power (i.e., accuracy over intermediate steps) on the graph domain. To achieve this goal, we pretrain a topological Schrodinger bridge matching (TSBM) model (yang2025topological) on cortical fMRI data and train SDENO match high-energy brain signals to their aligned low-energy states (van2013wu). The brain graph is constructed, followed by (glasser2016multi). The training of TSBM is conducted by using an alternating scheme and both forward and backward processes governed by the graph stochastic heat diffusion equation $dX_{t}=-cLX_{t}dt+g_{t}dW_{t}$, where $L$ is the normalized graph Laplacian. Fixing $X_{T}\sim p_{T}$ (prior) and $X_{0}\sim p_{0}$ (posterior), TSBM seeks to fit two (graph) neural networks $\phi_{\sharp}$ and $\phi_{\Box}$ for forward and backward SDEs for brain signals,
@@
[…]
Figure 6: Extrapolation performance on the OU process (left) and Heston model (right).
@@
We further explore the model’s performance under extrapolation settings. We consider two widely used SDEs, namely the Ornstein-Uhlenbeck (OU) process
$dX_{t}=-\theta(X_{t}-\mu)dt+\sigma dW_{t},$ (10)
and financial Heston model (de2018machine)
$dS_{t}=\mu S_{t}dt+\sqrt{V_{t}}S_{t}dW^{S}_{t},\,\,\,dV_{t}=\kappa(\theta-V_{t})dt+\zeta\sqrt{V_{t}}dW^{V}_{t},\,\,\,d\langle W^{S},W^{V}\rangle_{t}=\rho dt,$ (11)
where $S_{t}>0$ is the asset price and $V_{t}\geq 0$ is the instantaneous variance with $\sqrt{V_{t}}$ known as volatility, and $\rho\in[-1,1]$ represents the correlation between Brownian motions $W^{S}$ and $W^{V}$. The propagator for OU process is simply learned by one MLP whereas for Heston, after we simulate $W_{t}^{S}$, we obtain $W_{t}^{V}$ by letting $W_{t}^{V}=\rho dW^{S}_{t}+\sqrt{1-\rho^{2}}dW^{2}_{t}$, where $dW^{2}_{t}$ is another Wiener increment that is independent of $dW^{S}_{t}$. Then we concatenate their individual ($\xi^{S}$, $\xi^{V}$) and crossed Wick features ($\xi^{S}\xi^{V}$) and feed it to one additional MLP whose output is merged with estimated propagators by inner product. To achieve the extrapolation setting, we first sample 500 paths for each process with 100 steps with $t\in[0,1]$, and we feed our model the first 75 steps for training, whereas 100 steps for evaluation. The results are in Figure 6, one can see that our model not only fits the training data well but also demonstrates strong extrapolation power into the unseen future. This suggests a promising potential for our framework in forecasting applications.
5.3 Meta-Path, Meta-SDENO and Parameter Estimation
In this experiment, we consider a parametric family of SDEs with parameters $\boldsymbol{\gamma}\in\boldsymbol{\Gamma}$. We again consider the OU process in Eq. (10). We define Meta-paths as a family of trajectories generated by the OU process with different $\boldsymbol{\gamma}$ and Meta-SDENO denotes an operator model trained on these Meta-paths. Our hypothesis is if a Meta-SDENO shows good fit on Meta-paths, then it serves as a faithful forward model for parameter estimation under the ensemble Kalman filter (EnKF).
Stepwise Supervision
We let $\boldsymbol{\gamma}=[\theta,\mu,\sigma]\in\mathbb{R}^{3}$, and we set the ground truths to $\theta^{\star}=4$, $\mu^{\star}=1$ and $\sigma^{\star}=0.05$, respectively. With a fixed OU process structure, we initially sample $P=300$ paths with $T=300$, $dt=0.01$, and initial condition $X_{0}=0$. We train a Meta-SDENO model for 100 epochs using the same model architecture for the extrapolation experiment. More importantly, in this task, Meta-SDENO is trained using a stepwise supervision scheme. Specifically, within each epoch, we uniformly sample (then fix) $\ell=250$ combinations (priors) from predefined distributions, e.g., $\theta\sim U(0,5)$, $\mu\sim U(-3,3)$ and $\sigma\sim U(0,0.5)$. For each triplet, we reuse the same $dW$ channels obtained from the initial simulation process, producing 300 new paths (with shared driver) for this specific triplet. For time $t\in[0,\dots,T-1]$, we feed the model with the corresponding $dW$ records up to $t$ to let the model predict $\widehat{X}_{t+1}\in\mathbb{R}^{P}$. Finally, the total MSE loss of the model is $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{\ell\times P\times T}\sum_{l=1}^{\ell}\sum_{p=1}^{P}\sum_{t=1}^{T}\|\widehat{X}_{t+1}(p,l)-{X}_{t+1}(p,l)\|_{2}^{2}$.
Figure 7: Parameter estimation of the OU process.
EnKF and Parameter Estimation Process
After pretraining, we freeze Meta-SDENO and treat $\boldsymbol{\gamma}=[\theta,\mu,\sigma]\in\mathbb{R}^{3}$ as the latent state to be inferred with a standard EnKF (evensen2003ensemble; aanonsen2009ensemble). The EnKF iteratively adjusts the parameter ensemble so that model predictions match the observed Meta-paths, using the sample covariance between parameters and predicted observations to form a gain. This procedure reduces ensemble spread while steering the mean toward parameters that better explain the data. Full dynamics for EnKF is provided in Appendix F.6. We report the parameter estimate results as the average over the last five steps. As shown in Figure 7, after $300$ updates Meta-SDENO, coupled with EnKF, successfully recovers the ground truth parameters, suggesting a harmonic collaboration between Meta-SDENO and EnKF. We highlight that these findings could include further mean-field analyses of the EnKF within stochastic domains (law2016deterministic).
@@ 5.4 Manifold SDEs
Figure 8: Flood events prediction on $\mathbb{S}^{2}$.
Finally, we test our model on manifold SDEs. We pretrain the Riemannian Score-Based Generative Model (RSGM) (de2022riemannian) on the Flood dataset (Brakenridge2016). Let $X\in\mathbb{S}^{2}$, we consider the forward SDE as manifold Langevin dynamics, $dX_{t}=-\tfrac{1}{2}\nabla U(X_{t})\,dt+dW_{t}^{\mathcal{M}}$, where $\nabla$ is the Riemannian gradient. Setting $U(x)\equiv 0$ yields intrinsic Brownian motion on the compact manifold (i.e., $dX_{t}=dW_{t}^{\mathcal{M}}$), and the time-reversed diffusion becomes $dX_{t}=\nabla\log p_{T-t}(X_{t})\,dt+dW_{t}^{\mathcal{M}}$. After pretraining, we sample $dW^{\mathcal{M}}$ trajectories on the tangent bundle using the standard geodesic random walk (GRW) sampler (de2022riemannian). We then parallel-transport the recorded $dW^{\mathcal{M}}$ trajectory to a fixed reference plane to obtain Wick features. In this experiment, we build SDENO with two separate MLPs: one provides additional encoding on Wick features and the other approximates the ODE solutions. We apply a retraction (via the exponential map) to project both predictions and ground truth back onto $\mathbb{S}^{2}$. Finally, the model is trained via the MSE loss between trajectories on the tangent space. We note that RSGM is pre-trained with its original likelihood-based (score-matching) objective as in de2022riemannian, whereas our manifold SDENO is trained by minimizing an MSE loss on the tangent space. For evaluation, however, we place both models on equal footing by measuring the same NLL between their generated trajectories and the reference flood data, yielding $\mathrm{NLL}=0.49\pm 0.20$ for RSGM and $\mathrm{NLL}=0.51\pm 0.12$ for our manifold SDENO. Figure 8 shows a representative sample of SDENO generations on $\mathbb{S}^{2}$, and Table 3 summarises the NLL comparison with baseline methods. In Appendix F.7 (Figure 15), we visualize the prediction differences between SDENO and RSGM.

### v2

@@ 3.2 Chaos Expansion on SPDEs and SDEs
4 Model Architecture
$\mathcal{F}$-SPDENO
As discussed above, solutions of (1) admit a Wiener–chaos representation $X_{t}(x)=\sum_{\alpha\in\mathcal{J}}u_{\alpha}(t,x)\,\xi_{\alpha}$, such that, once the Q-Brownian motion $W$ and the initial condition $\chi_{0}$ are observed, the randomness is entirely carried by $\{\xi_{\alpha}\}$, while the propagators $u_{\alpha}(t,x)$ remain deterministic. In practice, we first compute the finite-dimensional Q-Brownian increments, then construct the Gaussian variables $\{\xi_{mj}\}$ to obtain the Wick features $\xi_{\alpha}$. These Wick features serve as the stochastic input to our backbone model, e.g., FNO. To make the propagators suitable to FNO targets, we further expand them in a temporal basis $\{\phi_{k}(t)\}_{k=1}^{K}$ as $u_{\alpha}(t,x)\approx\sum_{k=1}^{K}u_{\alpha,k}(x)\,\phi_{k}(t)$, so that we only need to learn the time-independent coefficient fields $\{u_{\alpha,k}(x)\}$ in FNO, and the final prediction is obtained by combining the FNO outputs with the chaos and temporal bases. Figure 2 summarises the whole work flow of $\mathcal{F}$-SPDENO.
SDENO
For the SDE case, we follow the same chaos-based representation and treat the propagators as deterministic functions of time that can be approximated by standard neural networks. As illustrated in the previous section, once the Brownian motion has been encoded into Wick features, each chaos coefficient $u_{\alpha}(t)$ satisfies a deterministic ODE in $t$, so one can use a wide range of architectures (even simple MLPs as in (Eigel and Miranda, 2024)) to approximate these time-dependent propagators. Concretely, we construct Wick features from the driving noise trajectory $\{W_{t}\}_{t\in[0,T]}$ and feed the time variable $t$ (or its positional encoding) into a backbone network, whose outputs are then combined with the Wick features via an explicit inner product. This yields a flexible SDENO family with different backbones: for example, we use a UNet for diffusion SDEs on images ($\mathcal{U}$-SDENO) and graph neural networks for diffusion interpolation on graphs ($\mathcal{G}$-SDENO). In $\mathcal{U}$-SDENO, the time dependence is modeled inside the UNet, whereas in $\mathcal{G}$-SDENO it is handled by a separate temporal network. See the experiment sections 5.2 for more details.
5 Numerical Experiments
In this section, we present a comprehensive empirical evaluation of our NO model across diverse settings: (i) solving SPDEs (the dynamic $\boldsymbol{\Phi}^{4}_{1}$ model and the 2D Navier-Stokes equation); (ii) solving SDEs in the image and graph domains, yielding a diffusion one step sampler and a topological interpolator; (iii) extrapolation on the financial Heston model; and (iv) learning SDE solution on manifold. All experiments are executed on an NVIDIA® H200 SXM GPU (141 GB HBM3e) within an HPC cluster. Additional experimental details and results, e.g., empirical parameter sensitivity, and parameter estimation with our generalized Meta-SDENO model, are provided in Appendix F.
5.1 Approximation of SPDE Solutions
Dynamic $\boldsymbol{\Phi}^{4}_{1}$
We begin to evaluate our $\mathcal{F}$-SPDENO model by the dynamic $\boldsymbol{\Phi}^{4}_{1}$ model, which is widely applied in the realm of superconductivity (Temam, 2012). The equation has the form
$dX_{t}=\bigl(\Delta X_{t}+3X_{t}-X_{t}^{3}\bigr)\,dt+dW_{t},$
$X(0,x)=\chi_{0}\in L^{2}(\mathbb{T}^{1}),\quad X_{t}(0)=X_{t}(1),$ (7)
where $(t,x)\in[0,T]\times[0,1]$, and $W$ is the Q-Brownian motion. We follow the experimental settings of Salvi et al. (2022); Hu et al. (2022) in generating sample paths, initial conditions, ground truth solutions, and train-test splits. We conduct two tasks with different training observations $N=1000$ and $10000$, respectively, and evaluate our model with two settings, (i) $W\mapsto X$ which assumes $W$ is observed but the initial condition is fixed all the time and (ii) $(X_{0},W)\mapsto X$ where $W$ is observed and $X_{0}$ changed across the samples.
We compare $\mathcal{F}$-SPDENO with various baselines such as NCDE, NCDE-FNO, NRDE (Lu et al., 2022b), FNO (Li et al., 2020) and NSPDE (Salvi et al., 2022). We note that we did not compare our model to architectures specifically designed for rough or even singular SPDEs, such as (Hu et al., 2022; Gong et al., 2023). In the one-dimensional dynamic $\boldsymbol{\Phi}^{4}_{1}$ setting considered here, the solution enjoys positive Hölder regularity (Li et al., 2025), so the regularity-structure machinery that underpins these models is not essential. The results are contained in Table 1. Our model achieves remarkable results relative to all other baselines, with lower $L^{2}$ errors across both small and large training schemes. Interestingly, increasing $N$ from $1000$ to $10000$ only leads to modest changes in the relative $L^{2}$ error for NSPDE and $\mathcal{F}$-SPDENO, especially on the $(X_{0},W)\mapsto X$ task. This behaviour is consistent with the fact that both methods already achieve very small errors in the $N=1000$ regime. In $\boldsymbol{\Phi}^{4}_{1}$, $N=1000$ trajectories already capture most of the variability of the dynamics, so increasing $N$ to $10000$ brings only limited additional benefit. This aligns with the recent benchmarking study in SPDE (Li et al., 2025).
Stochastic Navier-Stokes Equation
In the second experiment, we evaluate $\mathcal{F}$-SPDENO on the 2D incompressible Navier-Stokes equations in vorticity form (Li et al., 2020). The dynamic takes the form
$dX_{t}=\Bigl(\nu\,\Delta X_{t}\;-\;U[X_{t}]\!\cdot\!\nabla X_{t}\;+\;f\Bigr)\,dt+\sigma dW_{t},$ (8)
with $(t,x)\in[0,T]\times\mathbb{T}^{2},X_{0}=\chi_{0}\in L^{2}_{0}(\mathbb{T}^{2})$. Here we take $U[X_{t}]\coloneqq\nabla^{\perp}(-\Delta)^{-1}X_{t}$ with $\nabla^{\perp}\coloneqq(-\partial_{x_{2}},\partial_{x_{1}})$ where $\partial$ denotes the partial derivative of the spatial variable (i.e., $x=(x_{1},x_{2})$). The parameter $\nu>0$ is the viscosity; $f$ is a deterministic vorticity forcing. Similarly, we generate the samples and ground truth solutions following (Salvi et al., 2022). We train and evaluate our model on $\times64$ grid, and we also show the results on $\times16$ grid in Appendix F.2.
@@ 5.1 Approximation of SPDE Solutions
We present learning outcomes in Table 2 and Figure 3, following the qualitative protocol used in NSPDE (Salvi et al., 2022). $\mathcal{F}$-SPDENO attains the lowest error among all competitors in Table 2. Compared to the strongest baseline NSPDE, the error is reduced from $0.040$ to $0.037$ on $W\mapsto X$ (about $7.5\%$ relative improvement), and more substantially from $0.047$ to $0.031$ on $(X_{0},W)\mapsto X$ (about $34\%$ relative improvement). This indicates that our model better captures the joint effect of the random forcing and varying initial conditions. We also observe that deterministic operator learners such as FNO already perform competitively on $W\mapsto X$, but degrade notably when the initial condition varies (from $0.051$ to $0.073$). By contrast, $\mathcal{F}$-SPDENO consistently improves over FNO in both settings (from $0.051$ to $0.037$ and from $0.073$ to $0.031$), showing that enriching FNO with WCE-based noise features is crucial for handling stochastic SPDEs. In contrast to iterative inference methods like NSPDE, our framework first uses WCE to convert the stochastic forcing into static features. This enables a single FNO to learn the global solution operator in a one-shot forward pass, while still achieving superior accuracy.
Constrain on $\alpha$
In practice, the infinite chaos index set $\mathcal{J}$ in Eq. (4) is truncated by limiting the number of Brownian components ($I$), temporal basis modes ($J$), and the total Hermite order ($K$), yielding
$\mathcal{J}_{I,J,K}:=\Bigl\{\alpha=(\alpha_{ij})_{i,j\in\mathbb{N}}\in\mathbb{N}_{0}^{\mathbb{N}\times\mathbb{N}}:\ \alpha_{ij}=0\$
$\text{if }i>I\ \text{or }j>J,\ \ |\alpha|\leq K\Bigr\}\ \subset\ \mathcal{J},$
where $|\alpha|\coloneqq\sum_{i,j}\alpha_{ij}$. The condition $|\alpha|\coloneqq\sum_{i,j}\alpha_{ij}\leq K$ is crucial as it retains chaos indices that grow combinatorially (Neufeld and Schmocker, 2024). For non-singular dynamics such as $\boldsymbol{\Phi}^{4}_{1}$ and the 2D NS equation, we find that restricting to low-order $K\leq 2$ is sufficient in practice, without explicitly modeling higher-order information of the driving Brownian motion. In contrast, for singular SPDEs such as $\boldsymbol{\Phi}^{4}_{2}$ (Li et al., 2025), higher-order chaos components are expected to play an essential role. Since addressing such singular regimes requires additional analytical and modeling considerations, we leave the extension of WCE-based neural operators to this setting for future work.
[…]
Figure 6: Extrapolation performance on the Heston model.
@@ 5.2 Approximation on SDE Solutions
We now turn our attention to applying our model to SDEs. As mentioned, a key advantage of the NO method is that it enables one-shot evaluation. This naturally makes SDENO a one-step sampler for diffusion generative models (Ho et al., 2020; Song et al., 2020b). In this experiment, we apply our $\mathcal{U}$-SDENO on pretrained DDPM models on CIFAR10 and CelebA-HQ  [^2] Model id :google/ddpm-cifar10-32 and google/ddpm-celebahq-256.. Specifically, the canonical DDPM reverse SDE takes the form
$dX_{t}=\!\!\Bigl[-\tfrac{1}{2}\,\beta(t)\,X_{t}-\beta(t)\,\nabla_{x}\log p_{t}(X_{t})\Bigr]dt+\sqrt{\beta(t)}\,d\overline{W}_{t}.$
At each reverse step $t$, we sample $\widehat{X}_{t}$ from pretrained models, and compute the DDPM posterior mean $\mu_{t}$ and variance $\sigma_{t}^{2}$ from the fixed noise schedule $\{\beta_{s}\}_{s=1}^{T}$, and then draws $X_{t-1}=\mu_{t}+\sigma_{t}z_{t}$ with $z_{t}\sim\mathcal{N}(0,I)$ (no noise at t=0t{=}0). Operationally, $z_{t}$ realizes the reverse-time increment $d\overline{W}_{t}\approx\sqrt{\Delta t}\,z_{t}$. We sample $d\overline{W}$ trajectories followed by the previous work in (Zheng et al., 2023).
Figure 4 shows the generated images of $\mathcal{U}$-SDENO at the last prediction step compared with the pretrained DDPM outputs on CIFAR-10, and unconditional samples on CelebA-HQ. One can see that $\mathcal{U}$-SDENO reconstructs the images with high fidelity and produces unconditional samples of comparable quality, i.e., FID-50K of our model is 4.89 compared to the pretrained model 3.17 in CIFAR10. We provide a more detailed quantitative comparison between our model and other baselines in the Appendix F.3. Unlike standard diffusion models, which learn the score fields and integrate the reverse SDE step by step, SDENO learns a solution operator mapping a full noise trajectory $d\overline{W}_{1:T}$ directly to the final sample in a single forward pass. At sampling time, we only need to draw a Gaussian noise trajectory and evaluate SDENO once, without repeatedly invoking the large UNet denoiser. We finally note that the goal of this experiment is therefore not to compete with specialized one-step diffusion schemes, e.g., consistency models (Song et al., 2023), but to demonstrate that the same WCE-based NO framework can also handle high-dimensional SDEs in the image domain.
@@
In this experiment, we show SDENO also possesses strong interpolation power (i.e., accuracy over intermediate steps) on the graph domain. To achieve this goal, we pretrain a topological Schrodinger bridge matching (TSBM) model (Yang, 2025) on cortical fMRI data and train SDENO match high-energy brain signals to their aligned low-energy states (Van Essen et al., 2013). The brain graph is constructed, followed by (Glasser et al., 2016). The training of TSBM is proceeds by an alternating scheme, with the forward and backward processes governed by the graph stochastic heat diffusion equation $dX_{t}=-cLX_{t}dt+g_{t}dW_{t}$, where $L$ is the normalized graph Laplacian. Fixing $X_{T}\sim p_{T}$ (prior) and $X_{0}\sim p_{0}$ (posterior), TSBM seeks to fit two (graph) neural networks $\phi_{\sharp}$ and $\phi_{\Box}$ for forward and backward SDEs for brain signals,
@@
We further explore the model’s performance under extrapolation settings. We consider the popular financial Heston model (De Spiegeleer et al., 2018)
$dS_{t}=\mu S_{t}dt+\sqrt{V_{t}}S_{t}dW^{S}_{t},$
$dV_{t}=\kappa(\theta-V_{t})dt+\zeta\sqrt{V_{t}}dW^{V}_{t},\,\,\,d\langle W^{S},W^{V}\rangle_{t}=\rho dt,$
where $S_{t}>0$ is the asset price and $V_{t}\geq 0$ is the instantaneous variance with $\sqrt{V_{t}}$ known as volatility, and $\rho\in[-1,1]$ represents the correlation between Brownian motions $W^{S}$ and $W^{V}$. To learn the propagators for Heston, after we simulate $W_{t}^{S}$, we obtain $W_{t}^{V}$ by letting $W_{t}^{V}=\rho dW^{S}_{t}+\sqrt{1-\rho^{2}}dW^{2}_{t}$, where $dW^{2}_{t}$ is another Wiener increment that is independent of $dW^{S}_{t}$. Then we concatenate their individual ($\xi^{S}$, $\xi^{V}$) and crossed Wick features ($\xi^{S}\xi^{V}$) and feed it to one additional MLP whose output is merged with estimated propagators by inner product. To achieve the extrapolation setting, we first sample 500 paths with 100 steps with $t\in[0,1]$, and we feed our model the first 75 steps for training, whereas 100 steps for evaluation. The results are in Figure 6, one can see that our model not only fits the training data well but also demonstrates strong extrapolation power into the unseen future. This suggests a promising potential for SDENO in forecasting applications.
@@ 5.3 Manifold SDEs
Finally, we test our model on manifold SDEs. We pretrain the Riemannian Score-Based Generative Model (RSGM) (De Bortoli et al., 2022) on the Flood dataset (Brakenridge, 2016). Let $X\in\mathbb{S}^{2}$, we consider the forward SDE as manifold Langevin dynamics, $dX_{t}=-\tfrac{1}{2}\nabla U(X_{t})\,dt+dW_{t}^{\mathcal{M}}$, where $\nabla$ is the Riemannian gradient. Setting $U(x)\equiv 0$ yields intrinsic Brownian motion on the compact manifold (i.e., $dX_{t}=dW_{t}^{\mathcal{M}}$), and the time-reversed diffusion becomes $dX_{t}=\nabla\log p_{T-t}(X_{t})\,dt+dW_{t}^{\mathcal{M}}$. After pretraining, we sample $dW^{\mathcal{M}}$ trajectories on the tangent bundle using the standard geodesic random walk (GRW) sampler (De Bortoli et al., 2022). We then parallel-transport the recorded $dW^{\mathcal{M}}$ trajectory to a fixed reference plane to obtain Wick features. In this experiment, we build SDENO with two separate MLPs: one provides additional encoding on Wick features and the other approximates the ODE solutions. We apply a retraction (via the exponential map) to project both predictions and ground truth back onto $\mathbb{S}^{2}$. Finally, the model is trained via the MSE loss between trajectories on the tangent space. We note that RSGM is pre-trained with its original likelihood-based (score-matching) objective as in De Bortoli et al. (2022), whereas our manifold SDENO is trained by minimizing an MSE loss on the tangent space.
@@
For evaluation, however, we place both models on equal footing by measuring the same NLL between their generated trajectories and the reference flood data, yielding $\mathrm{NLL}=0.49\pm 0.20$ for RSGM and $\mathrm{NLL}=0.51\pm 0.12$ for our manifold SDENO.
Figure 7: Flood event prediction on $\mathbb{S}^{2}$.
Figure 7 shows a representative sample of SDENO generations on $\mathbb{S}^{2}$, and Table 3 summarises the NLL comparison with baseline methods. In Appendix F.7 (Figure 15), we visualize the prediction differences between SDENO and RSGM.